OpenStudy (anonymous):

4x - 6 - 8x - 9 = 21 Just checking again if I am right, But I got the answer.

OpenStudy (anonymous):

what's you answer?

OpenStudy (anonymous):

ughh.. I think 9 right ?

OpenStudy (anonymous):

negative 9.... :)

OpenStudy (anonymous):

you must have missed a negative sign somewhere...

OpenStudy (anonymous):

we can go through this if you want...

OpenStudy (anonymous):

oohh ryte I got it. from the negative four.

OpenStudy (anonymous):

good work... part of being good at math is checking your own work and see where your miscalculations occur... :)

OpenStudy (anonymous):

thank you. can I ask you one more question ?

OpenStudy (anonymous):

fire away... :)

OpenStudy (anonymous):

but when I put (4x-9) - 6 - (8x-9) - 9 = 1

OpenStudy (anonymous):

let's simplify this...

OpenStudy (anonymous):

(4x-9) - 6 - (8x-9) - 9 = 1 4x - 9 - 6 - 8x + 9 - 9 = 1 do you understand why the "-" changed to a "+" between the 8x and the 9?

OpenStudy (anonymous):

not rely...

OpenStudy (anonymous):

use distributive property...

OpenStudy (anonymous):